What are the domain and range for these equations?

  1. f(x)= square root of (x-1)
  2. f(x)=6x-1/1-2x
  3. the inverse of function f(x) = square root of (x-1)
1
Expert's answer
2021-05-07T10:05:49-0400

Solution.

1.

"f(x)=\sqrt{x-1}"


Domain: "x \in [1,\infty)."

Range: "y\in [0,\infty)."

2.


"f(x)=\frac{6x-1}{1-2x}"

Domain: "x \in R \backslash\{\frac{1}{2}\}."

"\lim_{x\to\pm\infty}\frac{6x-1}{1-2x}=\lim_{x\to\pm\infty}\frac{6-\frac{1}{x}}{\frac{1}{x}-2}=\frac{6}{-2}=-3."

So,

Range: "y\in (-\infty,\infty)\backslash\{-3\}."

3.


"f(x)=\sqrt{x-1},\newline
f^2(x)=x-1,\newline
x=f^2(x)+1,\newline
f^{-1}(x)=x^2+1, x\geq0."


Domain: "x\in [0,\infty)."

Range: "y\in [1,\infty)."